OpenStudy (anonymous):

Derivatives. y=(sec^-1 x) × √(x^2-1)

OpenStudy (kmeis002):

since you have two functions multiplying, this will be a product rule and a chain rule for the second function: \[\frac{d}{dx} \left [ arcsec(x) \sqrt{x^2-1} \right ] = \sqrt{x^2-1} \frac{d}{dx}\left [ arcsec(x) \right ] + arcsec(x) \frac{d}{dx} \left [ \sqrt{x^2-1} \right ]. \] Do you have an idea where to continue from there?

OpenStudy (anonymous):

yes, tnx, I was just making sure. That's what I have thought. when I do the product rule though, am I able to re-write sec^-1 x as x --------- sqrt(x^2-1) being that:  which would yield that\[y'=(\frac{x}{\sqrt{x^2-1}})(\frac{d}{dx}~\sqrt{x^2-1}~)~+~(\sqrt{x^2-1}~)(\frac{d}{dx}~\sec^{-1}x~) \] \[y'=(\frac{x}{\sqrt{x^2-1}})(\frac{1}{\sqrt{x^2-1}})~+~(\sqrt{x^2-1}~)(~\frac{x^2-1}{x})\]